OpenStudy (anonymous):

Silver (108 g) has a density 10.5g/cm^-3 the number of silver atom on a surface of area 10^-12 m^2 can be expressed in scientific notatation as y * 10^x the value x is?

OpenStudy (ghazi):

well calculate the number of atom first by using n=(mass)/molar mass...to find mass use mass= density*volume and already you have surface area.....so you can calculate volume

OpenStudy (anonymous):

V = 10 ^ -24 ???

OpenStudy (ghazi):

did you get this as volume?

OpenStudy (anonymous):

yes

OpenStudy (ghazi):

then find the mass of that surface area...after that number of atoms by multiplying number of moles with avagadro's number

OpenStudy (anonymous):

m = 10.5* 10^-18

OpenStudy (ghazi):

number of moles= mass/ 108

OpenStudy (anonymous):

lol...i just post to check..

OpenStudy (ghazi):

okay can you do it now?

OpenStudy (anonymous):

moles = 0.97 * 10^-19

OpenStudy (anonymous):

no of atoms = 0.97 * 6.02 * 10^23

OpenStudy (ghazi):

no

OpenStudy (ghazi):

you forgot 10^-19

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

atom = 5.83 * 10^4

OpenStudy (anonymous):

but the answer should be 7

OpenStudy (ghazi):

yes...i think so :)

OpenStudy (anonymous):

@Ishaan94 @IsraelYkb @Callisto @ghazi

OpenStudy (ghazi):

??

OpenStudy (anonymous):

the answer should be 7

OpenStudy (anonymous):

it shuld be 10^7

OpenStudy (anonymous):

@UnkleRhaukus

OpenStudy (anonymous):

@experimentX

OpenStudy (anonymous):

lol @ghazi

OpenStudy (ghazi):

@Yahoo! convert your density into gm/m^3 ..see area and density has different units...you'll have the answers.... kid :)

OpenStudy (anonymous):

i did...lol

OpenStudy (anonymous):

the ans given by me included..that

OpenStudy (anonymous):

@ghazi

OpenStudy (ghazi):

hmm...i understand lol

OpenStudy (anonymous):

V = 10 ^ -24 m^3 = 10 ^(-24+6) = 10^-18

OpenStudy (ghazi):

dude again mistake....volume= area*height ...how did you find height...and i can clearly state your question is mistaken :)

OpenStudy (anonymous):

i assumed it as Cube

OpenStudy (ghazi):

no...if you're assuming it as cube...then where's the height....secondly a surface area means to say in 2D you can't have thickness ...so how do you expect to calculate atoms??

OpenStudy (anonymous):

hm....hm.. u r correct..))

OpenStudy (anonymous):

so....is there a way to find the height

OpenStudy (anonymous):

@ghazi

OpenStudy (ghazi):

i guess NO

OpenStudy (anonymous):

LOL.... if it is not a cube we cant fine it

OpenStudy (ghazi):

nope...